A chemistry student needs to make 1000 ml of a 8% solution. How much 10% and 6% solutions need to be mixed?
Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(75887)

A chemistry student needs to make 1000 ml of a 8% solution. How much 10% and 6% solutions need to be mixed?
------
Equation:
active + active = active
0.10x + 0.06(1000-x) = 0.08*1000
10x + 6*1000 - 6x = 8*1000
4x = 2*1000
x = 500 ml (amt. of 10% solution needed)
---
1000-x = 500 (amt. of 6% solution needed)
==================
Cheers,
Stan H.

Answer by Alan3354(69443)

A chemistry student needs to make 1000 ml of a 8% solution. How much 10% and 6% solutions need to be mixed?
-------------
8 is the average of 10 & 6, so it's equal amounts.